-6 + 4/3(2) -1 =?
A. 1/3
B. -2/5
C. -1(1/3)
D. -5(1/5)
E. -6(1/3)

Answer by stanbon(75887)

-6 + (4/3)(2) -1 = ?
------------------
= -6 +(8/3) - 1
----
= -7 + 8/3
----
= (-21/3) + (8/3)
----
= -13/3
---
= -4 1/3
=================
Cheers,
Stan H.
=================
A. 1/3
B. -2/5
C. -1(1/3)
D. -5(1/5)
E. -6(1/3)